ULTRAHIGH-RESOLUTION SPECTROSCOPY OF NAPHTHALENE S_{1} ← S_{0} TRANSITION

\underline{SHUNJI KASAHARA}, KUMI AKITA, and MICHIRU YAMAWAKI , Molecular Photoscience Research Center, Kobe University, Nada-ku, Kobe 657-8501, Japan; MASAAKI BABA, Graduate School of Science, Kyoto University, Sakyou-ku, Kyoto 606-8502, Japan.

Several Doppler-free spectroscopic technique were applied to measure the rotational resolved high-resolution absorption and excitation spectra of the S_{1} ^{1}B_{1u} ← S_{0} ^{1}A_{g} electronic transition. Sub-Doppler excitation spectra and the Zeeman effects of the 0-0 ^{a} and several vibronic bands up to the excess energy around 2410 cm^{−1} were measured by crossing a laser beam perpendicular to a collimated molecular beam, and the Doppler-free absorption spectra of the bands of excess energy 435 cm^{−1} ^{b} and 1423 cm^{−1} were also measured by Doppler-free laser polarization labeling spectroscopy at room temperature. The molecular constants of the observed several vibrational states of the S_{1} ^{1}B_{1u}(v′) and S_{0} ^{1}A_{g}(v = 0) state were determined. The local energy shifts were found in the bands higher than 1423 cm^{−1} band, and identified as originating from the Colliolis interaction with other vibrational levels. The Zeeman splittings were found for all observed bands. The J, K-dependences of the observed splittings were studied. As a result, the observed Zeeman effects could be explained to be originating from the magnetic moment of the S_{1} ^{1}B_{1u} state induced by mixing with S_{2} ^{1}B_{3u} state by J-L coupling. ^{c}

aD. L. Joo, R. Takahashi, J. O’Reilly, H. Katô, and M. Baba, J. Mol. Spectrosc., 215, 155 (2002).

bM. H. Kabir, S. Kasahara, W. Demtröder, Y. Tatamitani, A. Doi, H. Katô, and M. Baba, J. Chem. Phys., 119, 3691 (2003).

cM. Okubo, J. Wang, M. Baba, M. Misono, S. Kasahara, and H. Katô, J. Chem. Phys., 122, 144303 (2005).